OpenStudy (anonymous):

Recall that linear approximation to the function f(x) at a is given by L(x)= f(a)+f'(a)(x-a) 1. Setting f(x)=x^4 and a=2 give an estimate of (1.999)4. 2. A quadratic approximation is given by Q(x)=f(a)+f'(a)(x-a)+1/2 f''(a)(x-a)^2. Use this formula to approximate (1.999)^4. 3. Use a calculator to determine the true value of (1.999)^4. Which approximation was better, the linear or quadratic? 4. Guess a formula for a cubic approximation.

OpenStudy (anonymous):

\(f'(x)=4x^3\) so \(f'(2)=32\)

OpenStudy (anonymous):

if \(a=2\) then \(f(a)=2^4=16\) and so \[L(x)=16+32(x-2)\] is the linear approximation

OpenStudy (anonymous):

if you put \(L(1.999)\) you get \[16+32\times (-.001)\]